# Toward a Physical Theory of Intelligence

### v1

@@ 2.3.2 Irreversibility and Conserved-Quantity Cost
Irreversibility arises when distinct encodings evolve into physically indistinguishable states, reducing the number of accessible encodings and forcing entropy to be exported into the environment through one or more conserved channels. In the bit-flip example of Appendix A, the encoding map $f:L\times C\to\mathcal{P}(P)$ becomes non-invertible whenever the control protocol $C_{t}$ deforms the metastable basins so that previously separated basins merge. During this interval the physical state no longer records which encoding the system originated from, and restoring this encoding requires exporting entropy associated with a conserved quantity [18, 21].
Encoding-level reversibility concerns whether the mapping from input to output encodings is bijective. By contrast, physical or energetic reversibility concerns whether the underlying physical evolution preserves all conserved quantities and produces no net entropy. Under the Conservation-Congruent Encoding (CCE) assumptions, encoding reversibility and conserved-quantity reversibility coincide [5, 6]; the reversible evolution of encoded information must track a reversible evolution of the relevant conserved measure.
When a set of encodings $\{\ell_{i}\}_{i=1}^{m}$ is merged into a single physically indistinguishable state, the minimal entropy export is determined by the Shannon entropy of the erased distribution:
@@
where $p_{i}$ is the prior probability of encoding $\ell_{i}$, and $\alpha$ denotes the fluctuation scale associated with the relevant conserved quantity (e.g., $k_{B}$ for thermal channels, $\hbar$ for angular-momentum channels, or more generally the scale that appears in the microscopic entropy measure of the reservoir). For equiprobable inputs this simplifies to
$S_{\min}^{\mathrm{CCE}}=\alpha\ln m.$ (15)
Let $\Delta S_{\min}^{(k)}$ denote the entropy exported during the $k$-th irreversible merge. If $N_{\mathrm{irrev}}(T)$ merges occur over the interval $[0,T]$, the cumulative minimal entropy production is
$S_{\min}^{\mathrm{CCE}}(T)=\sum_{k=1}^{N_{\mathrm{irrev}}(T)}\Delta S_{\min}^{(k)}.$ (16)
In the special case where the average Shannon entropy destroyed per merge is $H_{\mathrm{merge}}$, this admits the coarse-grained form
$S_{\min}^{\mathrm{CCE}}(T)=\alpha\,H_{\mathrm{merge}}\,N_{\mathrm{irrev}}(T),$ (17)
providing an intuitive “entropy-per-irreversible-operation” interpretation.
Energetic or conserved-quantity reversibility is achieved only when the encoding mapping remains bijective and the physical evolution maintains a one-to-one correspondence between encodings and the full set of conserved quantities relevant to the substrate. Whenever distinct encodings are merged, the resulting degeneracy must be resolved by exporting generalized free energy. In the general case, the minimal free-energy cost associated with erasing $m$ equiprobable encodings is
$\Delta F_{\min}^{\mathrm{CCE}}=\alpha\,\lambda\,\ln m,$ (18)
where $\lambda$ is the intensive variable conjugate to the conserved quantity controlling the reversible dynamics (e.g., temperature for thermal reservoirs, chemical potential for particle exchange, magnetic field for spin degrees of freedom, etc.). The classical Landauer bound [18, 6],
$\Delta F_{\min}=k_{B}T_{\mathrm{env}}\ln m,$ (19)
is recovered as the special case where the conserved quantity is internal energy and the conjugate variable is temperature.
2.4 GENERIC Formulation
The agent-environment model of §2.1 describes the joint process $(X_{t},E_{t})$ in discrete time through coupled transition kernels. To analyse how information is preserved or irreversibly lost within the agent, it is useful to introduce a continuous-time refinement of the agent’s internal dynamics. The GENERIC (General Equation for the Non-Equilibrium Reversible-Irreversible Coupling) formalism provides a standard description of nonequilibrium dynamics that enforces compatibility with conservation laws and the second law of thermodynamics [14, 20]. In the present setting, GENERIC describes the instantaneous reversible-dissipative flow underlying the coarse-grained updates $X_{t+1}\sim K_{X}(\cdot\mid X_{t},\Gamma_{E}(E_{t})).$ Under this refinement, the internal state $x_{t}\in X$ evolves according to a reversible-dissipative decomposition
$\dot{x}_{t}=J(x_{t},t)\,\nabla H(x_{t},t)-R(x_{t},t)\,\nabla\Xi(x_{t},t),$ (20)
where $J(x,t)$ is antisymmetric and generates a reversible, structure-preserving flow governed by a conserved quantity $H$, while $R(x,t)$ is symmetric positive semidefinite and generates dissipative relaxation through the dissipation potential $\Xi$. To ensure compatibility with nonequilibrium physics and conservation laws, we impose the standard GENERIC degeneracy conditions
$J(x,t)\,\nabla\Xi(x,t)=0,\qquad R(x,t)\,\nabla H(x,t)=0.$ (21)
These guarantee that the reversible flow preserves the dissipation potential $\Xi$, while the dissipative flow does not directly change the conserved quantity $H$. This cleanly separates the two physical modes of evolution. The instantaneous entropy-production rate associated with the dissipative component is
$\dot{S}_{\mathrm{prod}}(t)=\big\langle\nabla\Xi(x_{t},t),\;R(x_{t},t)\,\nabla\Xi(x_{t},t)\big\rangle\;\geq\;0,$ (22)
where the entropy measure is defined relative to the conserved quantity governing the reversible dynamics. Let $\alpha$ denote the fluctuation scale associated with this conserved quantity. We therefore define the corresponding irreversible-information rate as
$\dot{I}_{\mathrm{irr}}(t):=\dot{S}_{\mathrm{prod}}(t)/\alpha.$ (23)
Under the Conservation-Congruent Encoding (CCE) assumptions introduced in §2.3.1:
@@ 2.4 GENERIC Formulation
    the reversible component $J\nabla H$ preserves the metastable encoding structure, maintaining one-to-one correspondence between encodings and conserved-quantity configurations;
@@
    the dissipative component $R\nabla\Xi$ irreversibly destroys encodings by contracting phase-space volume along directions associated with conserved-quantity dissipation, corresponding to encoding merges and irreversible information loss under CCE.
This reversible-dissipative decomposition provides the dynamical backbone for subsequent sections. It underlies the interpretation of intelligence as a measure of useful work per unit irreversible information, the emergence of persistent informational structure associated with self-modelling, and the epistemic limits that follow from the non-invertibility of dissipative dynamics.
@@ 2.5 From Communication to Intelligence
Shannon’s communication theory established the fundamental limits on how much information can be transmitted across a physical channel subject to constraints such as noise and available signal power. Crucially, the theory treats communication as an information-preserving process: encodings are conveyed from sender to receiver, but no irreversible transformation of information need occur within the channel itself. Shannon’s bounds therefore quantify how efficiently a physical medium may transport information, not how efficiently a system may use information to guide its own evolution [25].
@@
Irreversible information processing is physically distinct from communication. Whenever encodings are collapsed, a physical system must export entropy associated with a conserved quantity. This generalized form of Landauer’s insight establishes that information use has intrinsic physical cost [18, 6]. The Conservation-Congruent Encoding (CCE) assumptions introduced in §2.3.1 ensure that such costs are well defined across all substrates.
@@
The present framework unifies these perspectives by relating communication to the irreversible use of information in performing goal-directed work. An intelligent system does not just receive or relay information; it irreversibly transforms information so that it constrains future behaviour and enables goal-directed physical work. Let $I_{\mathrm{irr}}$ denote the total irreversible information processed by the agent, and let $W_{\mathrm{goal}}$ denote the useful work performed on the environment along task-relevant degrees of freedom. We define the intelligence as
$\chi=\frac{W_{\mathrm{goal}}}{I_{\mathrm{irr}}},$ (24)
@@
representing the amount of goal-directed work extracted per nat of irreversible information processing. This ratio is well defined whenever $I_{\mathrm{irr}}>0$. In the limiting case $I_{\mathrm{irr}}=0$, the agent performs no irreversible transformations, preserves all encodings, and cannot generate any persistent, time-asymmetric influence on its environment. Such a fully reversible process therefore cannot be assigned a meaningful finite intelligence under this metric. Systems with larger $\chi$ convert their physically costly informational transformations into useful work more effectively.
This formulation extends Shannon’s analysis of communication and reversible information transport to intelligence and irreversible information processing [25]. It yields a substrate-independent, physically grounded measure of intelligence as the degree to which an agent converts irreversible information into useful, goal-aligned work.
@@
In time-dependent settings, intelligence admits a natural flux formulation. Let $\dot{I}_{\mathrm{irr}}(t)$ and $\dot{W}_{\mathrm{goal}}(t)$ denote the instantaneous rates of irreversible information processing and goal-directed work, respectively. We define the instantaneous intelligence as
$\chi(t)=\frac{\dot{W}_{\mathrm{goal}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (25)
@@
which characterizes the local efficiency with which irreversible information processing is converted into useful work at time $t$.
The physically meaningful measure of intelligence over a finite horizon $T$ is the ratio of integrated fluxes,
$\chi_{T}=\frac{\int_{0}^{T}\dot{W}_{\mathrm{goal}}(t)\,dt}{\int_{0}^{T}\dot{I}_{\mathrm{irr}}(t)\,dt},$ (26)
@@
which quantifies the total amount of goal-directed work extracted per nat of irreversibly processed information. Note that $\chi_{T}$ is defined as a ratio of integrated fluxes and does not, in general, equal the time integral of $\chi(t)$.
@@ 3 Quantitative Measures of Intelligence
Having established the coupled agent-environment dynamics, the representation of goals, and the physical definition of intelligence, we now outline three quantitative measures that characterize the informational and adaptive capacities of the system. These measures occupy distinct yet complementary roles. Information capacity describes the rate at which new information may enter through the agent’s observation ports. Identifiability capacity characterizes how sharply latent environmental structure is encoded in the agent’s trajectory. Intelligence quantifies how efficiently information is converted into goal-directed work. These measures draw on classical notions from communication theory, statistical estimation, and nonequilibrium physics, but are here unified within a single physical framework [25, 11, 29]. We conclude this section by deriving a hierarchy of formal bounds linking these quantities, showing how physical and informational constraints impose fundamental limits on a system’s achievable performance.
3.1 Information Capacity
The information capacity of the agent-environment coupling quantifies the local rate at which causal information about the environment enters the agent through its sensory ports. We define this quantity in terms of directed information, which captures causal information flow in systems with feedback [19, 27].
Let $\dot{I}_{E\rightarrow X}(t)$ denote the directed-information density under the projected interface measure $\bar{\mu}_{t}$. Writing $\Gamma_{E}(E^{t^{-}})$ and $\Gamma_{X}(X^{t^{-}})$ for the left limits of the environmental and internal port variables, respectively, we define
$\dot{I}_{E\rightarrow X}(t):=I_{\bar{\mu}_{t}}\!\big(\Gamma_{E}(E_{t});\,\Gamma_{X}(X_{t})\,\big|\,\Gamma_{E}(E^{t^{-}}),\Gamma_{X}(X^{t^{-}})\big),$ (27)
which measures the instantaneous causal information flow from the environmental port to the internal port.
Information capacity is thus a property of the agent-environment interface, rather than of the agent’s internal processing. The ports are the only variables through which environmental information can causally influence the agent; downstream transformation of this information is captured separately by the identifiability and intelligence measures. Finite sensing delays and internal dynamics are absorbed into the transition kernels and hence into the trajectory distribution $\bar{\mu}_{t}$.
The finite-horizon information capacity is defined as the temporal average of the directed-information density,
$\mathcal{C}_{T}[\bar{\mu}_{0:T}]:=\frac{1}{T}\int_{0}^{T}\dot{I}_{E\rightarrow X}(t)\,dt,$ (28)
representing the mean rate at which new environmental information becomes available to the agent’s internal dynamics over the interval $[0,T]$.
For interaction cycles of duration $\Delta t$, the continuous-time expression reduces to the per-step directed-information increment
$I_{E\rightarrow X}(t):=I_{\bar{\mu}_{t}}\!\big(\Gamma_{E}(E_{t});\,\Gamma_{X}(X_{t})\,\big|\,\Gamma_{E}(E^{t-1}),\Gamma_{X}(X^{t-1})\big),$ (29)
with finite-horizon information capacity
$\mathcal{C}_{T}[\bar{\mu}_{1:T}]:=\frac{1}{T}\sum_{t=1}^{T}I_{E\rightarrow X}(t),$ (30)
recovering the standard discrete formulation as $\Delta t\to 0$.
3.2 Identifiability Capacity (Fisher-Information Rate)
To quantify how sharply features of the agent-environment dynamics are encoded in the agent’s internal evolution, we introduce an identifiability capacity defined in terms of the Fisher-information rate [11, 29] of the induced path measure.
Let $\vartheta_{t}$ denote a (possibly time-varying) vector of parameters that index a family of transition densities $p_{\vartheta_{t}}(X_{t}\mid X_{[0,t^{-})})$ used to model the internal dynamics of the agent under the coupled agent-environment process. Variations in $\vartheta_{t}$ may represent latent environmental descriptors, slowly varying conditions, or abstract parameters of the agent’s dynamical model.
Define the instantaneous score function
$g_{t}(X_{[0,t]}):=\nabla_{\vartheta_{t}}\log p_{\vartheta_{t}}(X_{t}\mid X_{[0,t^{-}]}),$ (31)
which measures the local sensitivity of the trajectory likelihood to perturbations of $\vartheta_{t}$. The corresponding continuous-time Fisher-information rate is
$\mathcal{F}_{T}[\mu_{0:T}]=\frac{1}{T}\int_{0}^{T}\mathrm{tr}\,\mathbb{E}_{\mu_{t}}\!\left[g_{t}(X_{[0,t]})\,g_{t}^{\top}(X_{[0,t]})\right]dt.$ (32)
This quantity measures the average curvature of the log-likelihood of the agent’s trajectory with respect to $\vartheta_{t}$ and therefore quantifies how precisely variations in the model parameters are reflected in the observed internal evolution.
For interaction cycles of duration $\Delta t$, the continuous-time expression reduces to the discrete-time score
$g_{t}(X^{t+1})=\nabla_{\vartheta_{t}}\log p_{\vartheta_{t}}(X_{t+1}\mid X^{t}),$ (33)
with Fisher-information rate
$\mathcal{F}_{T}[\mu_{1:T}]=\frac{1}{T}\sum_{t=1}^{T}\mathrm{tr}\,\mathbb{E}_{\mu_{t}}\!\left[g_{t}(X^{t+1})\,g_{t}^{\top}(X^{t+1})\right],$ (34)
recovering the discrete formulation as $\Delta t\to 0$.

### v2

@@ 2.3.3 Irreversibility and Conserved-Quantity Cost
Within the quotient-based framework of §2.3.1, irreversible information loss corresponds to a many-to-one mapping between dynamically induced equivalence classes defined by statistically indistinguishable macroscopic trajectories under a chosen multiplexed coarse-graining. Such a collapse of equivalence classes necessarily entails the destruction of physically realizable distinctions and therefore requires entropy export to degrees of freedom not retained by the system. This establishes a general connection between irreversibility and entropy production that is independent of any particular encoding scheme.
The Conservation-Congruent Encoding (CCE) conditions introduced in §2.3.2 identify a broad class of systems for which this entropy cost can be made explicit. Under CCE, abstract equivalence classes admit metastable physical realizations, and irreversible information loss occurs when control protocols deform the metastable basin structure so that previously distinct basins merge. During such operations, the physical state no longer encodes which equivalence class the system originated from, and restoring this information necessarily requires exporting entropy via the specific conserved quantities governing the dissipative flow to the environment [15, 17].
Encoding-level reversibility concerns whether the induced mapping between equivalence classes remains bijective at the operational level of description. Conserved-quantity reversibility, by contrast, concerns whether the underlying physical evolution preserves the relevant conserved measures and produces no net entropy. Under the CCE assumptions, these notions coincide: reversible manipulation of encoded information must correspond to reversible evolution with respect to the conserved quantities of the environmental baths coupled to the operational channels [4, 5].
Consider an irreversible operation in which a set of equivalence classes $\{\ell_{i}\}_{i=1}^{m}$ is merged into a single physically indistinguishable state. Under the CCE assumptions, the minimal entropy export required to perform this operation is bounded below by the Shannon entropy of the erased distribution,
@@
where $p_{i}$ denotes the prior probability of class $\ell_{i}$, and $\alpha$ is the characteristic entropy scale associated with the conserved quantity (e.g., $k_{B}$ for thermal channels or $\hbar$ for angular momentum). For equiprobable inputs this bound reduces to $S_{\min}^{\mathrm{CCE}}=\alpha\ln m$, recovering Landauer-type limits as a special case [15, 5].
While the metrics developed in §2.3 fundamentally compare macroscopic goal-directed influence to abstract information-processing costs, it is often instructive to evaluate them in a common physical currency. When we seek to frame these measures explicitly as dimensionless physical efficiencies (e.g., Joules of structural work per Joule of computational cost), we require an energetic conversion. This is naturally provided by standard conjugacy relations: the irreversible export of a conserved quantity across a boundary occurs against an intensive conjugate force $\mathcal{F}$ fixed by the environment. Multiplying the scaled informational erasure by this conjugate force yields the minimal energetic cost of the computation, allowing for direct comparison against the macroscopic physical footprint [7].
Under CCE, the Shannon term in $S_{\min}^{\mathrm{CCE}}$ is precisely the dimensionless irreversibility of the induced many-to-one map on equivalence classes, while the prefactor sets the physical scale of the associated export. The corresponding generalised energetic Landauer bound is
$E_{\min}^{\mathrm{CCE}}$ $=\mathcal{F}\,S_{\min}^{\mathrm{CCE}}=-\big(\mathcal{F}\alpha\big)\sum_{i=1}^{m}p_{i}\ln p_{i}.$ (26)
This makes explicit the bridge from abstract irreversibility to energetic cost: the trajectory-level asymmetry measures of §2.3.1 supply the dimensionless log-likelihood ratio (hence the Shannon term), and CCE supplies the conversion factor $(\mathcal{F}\alpha)$ into Joules.
For thermal entropy export, $\mathcal{F}=T$ and $\alpha=k_{B}$, yielding the usual Landauer form
$E_{\min}$ $=k_{B}T\,\Big(-\sum_{i=1}^{m}p_{i}\ln p_{i}\Big).$ (27)
For an encoding stabilised by angular momentum, $\mathcal{F}=\omega$ and $\alpha=\hbar$, giving
$E_{\min}$ $=\hbar\omega\,\Big(-\sum_{i=1}^{m}p_{i}\ln p_{i}\Big).$ (28)
Other conserved channels admit analogous expressions with their corresponding conjugate forces.
Real agents typically flux multiple conserved quantities. If an irreversible operation exports entropy through channels indexed by $a$, each with characteristic entropy scale $\alpha_{a}$ and conjugate force $\mathcal{F}_{a}$, then the unified energetic lower bound is additive across channels,
$E_{\min}^{\mathrm{CCE}}$ $=\sum_{a}\mathcal{F}_{a}\,S_{\min,a}^{\mathrm{CCE}}=-\Big(\sum_{a}\mathcal{F}_{a}\alpha_{a}\Big)\sum_{i=1}^{m}p_{i}\ln p_{i},$ (29)
where the final form applies when the same erased distribution $\{p_{i}\}$ is implemented across channels under CCE. Interpreting $\boldsymbol{\alpha}=(\alpha_{a})_{a}$ and $\mathbf{F}=(\mathcal{F}_{a})_{a}$, the prefactor is the inner product $\mathbf{F}\cdot\boldsymbol{\alpha}$.
In cases where the chosen operational interface fails to align with a metastable encoding, the irreversible collapse of dynamically induced equivalence classes still necessitates physical export. Crucially, however, this lack of a clean encoding is an epistemic artifact of the measurement, not a physical breakdown. Because the chosen coarse-graining does not adequately track the relevant conserved quantities, the physical costs leak into unmonitored degrees of freedom. Consequently, while the fundamental conservation laws remain exact, the observed cost becomes highly distributed and evades a clean, closed-form expression at that specific level of description.
2.4 System Boundaries and Structured Information Processing
The information-processing framework developed in §2.3 characterises irreversible computation in terms of path-level distinguishability under a chosen multiplexed operational coarse-graining $\Phi$. Such quantities are well defined only once the physical limits of the system have been specified. In this section, we formalise the notion of a system boundary in a manner consistent with the joint agent–environment dynamics of §2.1 and the path-space framework of §2.3, treating both the boundary and the internal encodings as derived, scale-dependent objects rather than primitive assumptions.
Throughout, the base joint dynamics of the universe $(X,E)$ are assumed to be strictly microscopic and deterministic. The stochastic process $(X_{t},E_{t})$ evolving under a joint transition kernel, as introduced in §2.1, strictly emerges as a mathematical consequence of marginalising over unobserved degrees of freedom. We now treat the choice of these boundary variables—alongside the choice of internal operational variables—as a rigorous joint search problem driven by the optimization of physical efficiencies.
We take $X$ to denote the candidate microscopic system variables. A fundamental requirement is that $X$ itself forms a dynamically connected system. At a given multiplexed coarse-graining $\Phi:X\rightarrow Z_{1}\times\dots\times Z_{C}$, this means that the internal degrees of freedom cannot be decomposed into a nontrivial partition whose components evolve independently while preserving the same path-level informational and thermodynamic structure. This requirement excludes purely abstract collections of variables and ensures that $X$ corresponds to a single cohesive physical system rather than a disjoint aggregate.
Given a connected system $X$, defining the agent as an open thermodynamic entity requires specifying two distinct mappings: the physical boundary projections $(\Pi_{X},\Pi_{E})$ that mediate environmental interaction, and the internal multiplexed projection $\Phi$ that defines the operational state. These maps are not assumed a priori. Instead, they are jointly selected so that the induced internal dynamics satisfy a strict set of criteria, ultimately serving to optimise the system’s operational goals:
@@ 2.4 System Boundaries and Structured Information Processing
    Dynamical Closure and Noise Generation: Environmental influence on the internal evolution must enter only through the boundary variables. Because the boundary projections $\Pi_{X}:X\rightarrow\Xi_{X}$ and $\Pi_{E}:E\rightarrow\Xi_{E}$ map the deterministic microscopic reality down to lower-dimensional interface ports, they strictly define the unresolved degrees of freedom that act as the environmental bath. Variations in microscopic environmental degrees of freedom orthogonal to $\Pi_{E}(E_{t})$ are marginalised, generating the stochastic noise floor of the induced transition kernel $K_{X}$. The boundary must be selectively permeable, constraining exactly which physical fluxes (and thus which conjugate forces) are allowed to perturb the system.
@@
    Encoding Preservation (Maximising the Numerator): Beyond closure, admissibility requires the preservation of structured information processing against the boundary-induced noise. The internal operational projection $\Phi$ must map to a composite quotient space $Z$ that aligns with the system’s Conservation-Congruent Encodings (CCE). The joint selection of the interface triplet $(\Pi_{X},\Pi_{E},\Phi)$ must preserve the multidimensional metastable basins required for meaningful prediction and control. Coarsening the boundary or the internal mapping too aggressively collapses these task-relevant macroscopic distinctions, directly reducing the macroscopic goal-directed work $W_{\mathrm{causal},T}$ that the system can perform.
  3. (iii)
    Minimality (Minimising the Denominator): Among closed and preserving architectures, the optimal interface $(\Pi_{X},\Pi_{E},\Phi)$ is selected by minimising unnecessary physical cost. Boundaries that expose superfluous internal variables ($\Pi_{X}$) or track irrelevant environmental fluctuations ($\Pi_{E}$) force the system to process and eventually erase non-predictive noise. This, along with internal encodings ($\Phi$) that track irrelevant physical channels, introduces unneeded physical dissipation. It inflates the system’s irreversible computational cost, captured by the generalised multi-channel Landauer sum $E_{\min}^{\mathrm{CCE}}=\sum_{c}\mathcal{F}_{c}\alpha_{c}S_{\min}^{(c)}$, without improving prediction or control, thereby degrading its overall efficiency.
Under this definition, identifying a macroscopic agent becomes a formal optimization problem. We treat the identification of the system and its computational states as a fundamental modelling decision: the most physically meaningful description of an agent is precisely the joint choice of boundary $(\Pi_{X},\Pi_{E})$ and internal multiplexed encoding $\Phi$ that simultaneously maximises its intelligence ($\chi$) and consciousness ($\kappa$) metrics. Once this optimal interface is identified, the stochastic port-based formulation of §2.1 and the path-space thermodynamics of §2.3 click together to provide a faithful, energetically bounded representation of the agent.
@@ 2.5 From Communication to Intelligence
Shannon’s communication theory established the fundamental limits on how much information can be transmitted across a physical channel subject to constraints such as noise and available signal power. Crucially, the theory treats communication as a process whose objective is faithful transmission rather than irreversible transformation. Shannon’s bounds therefore quantify how efficiently a physical medium may transport information, not how efficiently a system may use information to guide its own evolution [22].
@@
Irreversible information processing is physically distinct from communication. Whenever encodings are collapsed, a physical system must export entropy associated with a conserved quantity. This generalised form of Landauer’s insight establishes that information use has intrinsic physical cost [15, 5]. The Conservation-Congruent Encoding (CCE) assumptions introduced in §2.3.2 ensure that such costs are well defined across all substrates.
@@
The present framework unifies these perspectives by relating communication to the irreversible use of information in performing goal-directed work. An intelligent system does not just receive or relay information; it irreversibly transforms information so that it constrains future behaviour and enables goal-directed physical work. Let $W_{\mathrm{causal}}$ denote the useful work performed on the environment (as derived via the spatiotemporal boundary value problem in §2.2), and let $I_{\mathrm{irr}}$ denote the total irreversible information processed by the agent (bounded by the CCE limits in §2.3.3). We define the intelligence as
$\chi=\frac{W_{\mathrm{causal}}}{I_{\mathrm{irr}}},$ (30)
@@
representing the amount of goal-directed work extracted per nat of irreversible information processing. This ratio is well defined whenever $I_{\mathrm{irr}}>0$. In the limiting case $I_{\mathrm{irr}}=0$, the agent performs no irreversible transformations, preserves all encodings, and cannot generate any persistent, time-asymmetric influence on its environment. Such a fully reversible process therefore cannot be assigned a meaningful finite intelligence under this metric. Systems with larger $\chi$ convert their physically costly informational transformations into useful work more effectively. This definition does not measure task success or reward accumulation directly, but rather the physical efficiency with which irreversible informational transformations are converted into work aligned with externally specified goals.
@@
In time-dependent settings, intelligence admits a natural flux formulation. Let $\dot{I}_{\mathrm{irr}}(t)$ and $\dot{W}_{\mathrm{causal}}(t)$ denote the instantaneous rates of irreversible information processing and work, respectively. We define the instantaneous intelligence as
$\chi(t)=\frac{\dot{W}_{\mathrm{causal}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (31)
@@
which characterises the local efficiency with which irreversible information processing is converted into useful work at time $t$. The physically meaningful measure of intelligence over a finite horizon $T$ is the ratio of integrated fluxes,
$\chi_{T}=\frac{\int_{0}^{T}\dot{W}_{\mathrm{causal}}(t)\,dt}{\int_{0}^{T}\dot{I}_{\mathrm{irr}}(t)\,dt},$ (32)
@@
which quantifies the total amount of goal-directed work extracted per nat of irreversibly processed information. Note that $\chi_{T}$ is defined as a ratio of integrated fluxes and does not, in general, equal the time integral of $\chi(t)$. This reflects the fact that intelligence is not an instantaneous property of a system, but an emergent efficiency defined over irreversible transformations accumulated across time.
@@ 3 Quantitative Measures of Intelligence
Having established the coupled agent–environment dynamics, the representation of goals, and the notion of admissible system boundaries, we now formalise the core quantitative measures of the present framework. These measures characterise intelligence as a physical efficiency arising from irreversible and preserved information processing, rather than as a behavioural or algorithmic property.
We introduce two complementary quantities. Intelligence quantifies how efficiently irreversible information processing is converted into goal-directed physical work. Consciousness quantifies how efficiently preserved, structure-maintaining information supports the same process, capturing the role of persistent internal organization in enabling sustained, long-horizon behaviour. Together, these quantities separate the contributions of irreversible intervention and reversible structural reuse in physically embodied intelligent systems.
Both measures are defined directly in terms of physically measurable quantities—goal-directed work and irreversibly or reversibly processed information—and are evaluated relative to a chosen admissible system boundary and operational coarse-graining, as defined in §2.4.